# PointMapPolicy: Structured Point Cloud Processing for Multi-Modal Imitation Learning

### v1

@@ 1 Introduction
The advent of diffusion-based Imitation Learning (IL) has allowed robots to carry out complex, long-horizon tasks from raw image observations [1, 2]. RGB images are a common observation modality for diffusion policies due to their ubiquitousness and rich semantic information. However, policies conditioned on only RGB images lack 3D geometric information about the scene. This 3D information is crucial for learning generalizable policies that can act precisely in complex 3D scenes, especially when using multiple camera views [3, 4, 5, 6]. An alternative modality is point clouds, unstructured sets of 3D points that preserve geometric shape, distances, and spatial relationships. In addition, points captured from multiple camera views can be transformed into a common reference frame and concatenated, yielding a natural and powerful way to fuse multiple cameras. Although numerous works use point clouds as an input modality [7, 8, 9], their irregular structure limits the network architectures that can be used with them. In contrast, RGB images are on a regular grid and can be processed using convolutional operators, but are susceptible to changes in perspective and lighting.
@@
In this paper, we take inspiration from recent advances from the computer vision community in stereo reconstruction [10, 11] to propose using point maps, an observation modality that is novel in diffusion-based imitation learning. Point maps encode 3D information in a regular, 2D grid of Cartesian coordinates. This results in a structured data type that can be used with standard architectures such as ResNet [12], ViT [13], or ConvNeXt [14]. This obviates the need for steps like K-Nearest Neighors (KNN) and Farthest Point Sampling (FPS) [15], which are computationally expensive operations common to point cloud methods [16, 17, 6]. At the same time, because they are geometrically grounded, point maps from multiple views can be transformed into the same reference frame, increasing robustness to perturbations in camera perspective.
@@
We integrate point maps into a standard diffusion-based imitation learning framework based on EDM [18] to demonstrate their effectiveness as a drop-in replacement for RGB images or point clouds. We validate the effectiveness of point map observations on two challenging benchmarks: RoboCasa [19] and CALVIN [20]. These benchmarks feature language-conditioned tasks and diverse scenes, and require spatial reasoning and long-term planning. Across both benchmarks, point map-based policies outperform baselines using RGB, depth maps, or point clouds, demonstrating superior learning efficiency and generalization. Our method is computationally efficient in training and inference, sometimes by an order of magnitude.
@@
Contributions: Our contributions are the following: 1) we propose PointMapPolicy (PMP), a method for diffusion-based imitation learning on point maps, a powerful observation modality that has never been used in diffusion imitation learning; 2) we achieve state of the art results among policies trained from scratch on the CALVIN benchmark [20], and outperform other observation modalities on RoboCasa [19]; 3) we present systematic ablations of vision backbones (e.g. ResNet [12], ViT [13], ConvNeXt [14]) and paradigms for fusing color and geometries information. Our source code will be made publicly available upon publication.
@@ 2 Related Work
2D Visual Representations for Imitation Learning. Recent imitation learning approaches [1, 21, 22, 23, 24] rely predominantly on 2D visual representations such as RGB images or videos. Such representations are widely utilized due to their capacity to capture rich textural and semantic information, as well as their accessibility through low-cost cameras. However, 2D image modalities have inherent limitations: they contain 3D information only implicitly, are vulnerable to viewpoint and lighting changes and occlusions, and typically underperform in tasks requiring detailed spatial reasoning and geometric alignment [3, 4, 5, 6].
@@
3D Visual Representations for Imitation Learning. To overcome these limitations, a growing amount of research incorporates explictly 3D representations such as depth maps, point clouds, or voxels. Voxel-based methods like C2F-ARM [25] and Perceiver-Actor [26] voxelize point clouds and use a 3D-convolutional network for action prediction, but require high voxel resolution for precision tasks, resulting in high memory consumption and slow training. DP3 [7] encodes sparse point clouds using FPS, followed by a lightweight MLP to produce a compact embedding vector of the observation. While efficient, this approach discards local geometric structure that can be critical for fine-grained tasks. In contrast, 3D Diffuser Actor [27] computes tokens by lifting 2D image features into 3D space by using averaged depth information and camera parameters, and applies FPS after the first cross-attention layer. FPV-Net [28] fuses RGB and point cloud modalities by injecting global and local image features into a point cloud encoder using adaptive normalization layers, but is still limited by the disadvantages of both modalities.
@@
Multi-View Representation. Complementary work, such as Robot Vision Transformer (RVT) [29], avoids working directly with raw point clouds by proposing a novel multi-view representation. This approach re-renders the point cloud from a set of orthographic virtual cameras, deriving a 7-channel point map (RGBD + XYZ) from each view. RVT-2 [30] improves this approach for high-precision tasks by introducing a multi-stage inference pipeline: it first identifies a region-of-interest, truncates the observation to this area of interest, and then runs policy inference. However, neither of these methods use action diffusion, instead relying on key-frame based manipulation with a motion planner [31]. Furthermore, geometric and color information are fused naively at the channel level, whereas we investigate more sophisticated techniques for fusion.
@@
Diffusion-Policy Backbones. Due to the non-Markovian nature of human demonstrations, where successful decision-making often depends on histories of past observations and actions, early work used RNN-based architectures [32], but struggled with vanishing gradients and limited scalability. This led to the adoption of Transformer-based architectures, which offer global attention and parallelism, enabling superior performance in tasks requiring long-horizon reasoning [33, 34, 35], becoming the standard backbone for many methods [7, 27, 28, 21]. However, Transformers are computationally intensive and scale quadratically with the sequence length, which limits the number of tokens that can be used to encode the observation.
@@
To mitigate these challenges, recent works [36, 37] explore State Space Models (SSMs) like Mamba [38], achieving linear-time complexity and improved sample efficiency, particularly in low-data regimes. Additionally, recent recurrent architectures such as xLSTM [39] provide an appealing balance, maintaining the temporal modeling strengths of traditional RNNs while introducing architectural innovations that improve gradient flow and expressiveness. Despite being less expressive than self-attention, xLSTM significantly reduces compute and memory costs, making it well-suited for real-time or resource-constrained applications. X-IL [37] systematically compares different architectural parts, and finds that xLSTM performs competitively with Transformers in multi-modal imitation learning. Building on these insights, PointMapPolicy adopts xLSTM as its diffusion backbone, balancing temporal modeling capability with efficient training and inference.
@@ 3.1 Problem Formulation
Imitation Learning aims to learn a policy from expert demonstrations. Given a dataset of expert trajectories $\mathcal{D}_{\mathbf{\tau}}=\{\mathbf{\tau}_{i}\}_{i=1}^{N}$, where each trajectory $\mathbf{\tau}_{i}=\left((\mathbf{s}_{1},\mathbf{a}_{1}),(\mathbf{s}_{2},\mathbf{a}_{2}),\dots,(\mathbf{s}_{K},\mathbf{a}_{K})\right)$. The objective is to learn a policy $\pi(\bm{\bar{a}}|\mathbf{s})$ that maps observations $\mathbf{s}$ to a sequence of actions $\bm{\bar{a}}=(\mathbf{a}_{k},\mathbf{a}_{k+1},\dots,\mathbf{a}_{k+H})$. Predicting sequences of actions, i.e. action chunking, allows for more temporally consistent action prediction [40]. Each observation $\mathbf{s}$ contains multi-view RGB-D images and language instruction for the current trajectory.
@@ 3.2 Score-based Diffusion Policy
Our approach employs the EDM framework for continuous-time action diffusion [18, 41] to generate actions. Diffusion models are generative models that learn to generate new samples through learning to reverse a Gaussian Perturbation process. In PointMapPolicy, we apply a score-based diffusion model to formulate the policy representation $\pi_{\theta}(\bm{\bar{a}}|\mathbf{s})$. This perturbation and its inverse process can be expressed through a Stochastic Differential Equation (SDE):
@@
After training, we can generate new action sequences beginning with Gaussian noise by iteratively denoising the action sequence with a numerical Ordinary Differential Equation (ODE) solver. Our approach utilizes the DDIM-solver, a specialized numerical ODE-solver for diffusion models [43] that enables efficient action denoising in a minimal number of steps. Across all experiments, our method uses $4$ denoising steps.
@@ 3.3 Observation Tokenization
We are given an observation $\mathbf{s}_{k}$ in step $k$ as well as a textual language instruction $z_{\text{lang}}$. The language instruction is first tokenized using a pretrained CLIP text encoder [44] to generate language embeddings. For RGB inputs, we use Film-ResNet [45] with pretrained ImageNet weights to generate visual embeddings from the observation $\mathbf{s}_{k}$.
@@ 3.4 PointMapPolicy
Backbones. Given the multi-modal tokens from Section 3.3, a learnable positional embedding is added to each token. PMP uses a decoder-only backbone from X-IL [37] with x-LSTM as the core computational unit. All tokens are concatenated as inputs to the X-Block, which is the diffusion score network $D_{\theta}$. While Transformers dominate most imitation learning policies, X-IL demonstrated that the recent recurrent architecture xLSTM excels in robot learning tasks. The core computational element within X-Block is the m-LSTM layer, which serves an analogous function to self-attention in Transformer architectures. The denoised action tokens produced by the X-Block are then used to guide the robot’s behavior, resulting in a policy that effectively leverages both the geometric precision of point maps and the rich semantic understanding from RGB images.
@@ 4 Simulation Experiments
Figure 4: Overview of Simulation and Real World Experiments used to test PointMapPolicy. From left to right: CALVIN Benchmark [20], RoboCasa [19], and our Real World Setup.
We conduct experiments on two simulation benchmarks RoboCasa [19] and CALVIN [20]. We aim to answer the following questions: Q1) How does PointMapPolicy compare to state-of-the-art 2D and 3D imitation learning policies? Q2) How do the fusion methods perform compared to other modalities? Q3) Can current vision encoders effectively extract the geometric and semantic information from point maps required for robust decision-making?
RoboCasa: The RoboCasa benchmark [19] is a large-scale simulation framework designed to evaluate IL agents across a wide range of household manipulation tasks. Built on a physically realistic environment with rich visual rendering, RoboCasa supports task diversity, long-horizon behaviors, and fine-grained physical interactions, making it a compelling testbed for assessing both generalization and behavior diversity in policy learning. We use the RoboCasa benchmark to assess whether our proposed point map representation can enable effective learning and generalization across manipulation tasks of increasing complexity, object count, and behavioral variation.
CALVIN: The CALVIN benchmark [20] provides a large-scale framework for evaluating language-conditioned IL policies in visually rich, long-horizon manipulation tasks. The benchmark contains 34 distinct manipulation tasks such as button-pressing, drawer-opening, object-picking, and pushing. Each rollout consists of a sequence of 5 language instructions, and the agent must complete one task before proceeding to the next. Policies are evaluated on 1,000 such instruction chains per seed, and success is measured by the average number of correctly completed tasks in each sequence.
Baselines: For RoboCasa, we benchmark against Behavioral Cloning (BC) [19], GR00T-N1 [46], 3D Diffusion Policy (DP3) [7], and 3D Diffuser Actor (3DA) [27]. Note that GR00T-N1 results use $100$ demonstrations, while all other methods use $50$ human demonstrations. To systematically evaluate the effectiveness of our representation, we further compare other against image-based baselines using only RGB data (RGB), and only depth data (Depth). We then compare PMP against multiple variants introduced in Section 3.4: PMP-6ch directly uses 6-channel point maps as inputs, and PMP-xyz only uses xyz coordinates as inputs. All five methods share the same architectures and parameters for fair comparison. Details can be found in Appendix 6.
For CALVIN, we primarily compare our approach against models without robot-specific pretraining, though we include all results for reference. DP3, 3DA, and CLOVER [47] are selected as representative policies using RGB-D inputs. MDT [48], MoDE [23], and Seer [49] are selected as RGB-based policies. Further details of these baselines can be found in Appendix A.2.
@@
Table 1: Success rate (%) for each task in RoboCasa [19]. The models were trained for 50 epochs with 50 human demonstrations per task and evaluated with 50 episodes for each task. The bold numbers highlight the best achieved success rate for that task among all the models.
@@
SuSIE [51] $\checkmark$ Diffusion 87.0% 69.0% 49.0% 38.0% 26.0% 2.69
GR-1 [52] $\checkmark$ Cont. 85.4% 71.2% 59.6% 49.7% 40.1% 3.06
OpenVLA [22] $\checkmark$ Discrete 91.3% 77.8% 62.0% 52.1% 43.5% 3.27
RoboDual [53] $\checkmark$ Diffusion 94.4% 82.7% 72.1% 62.4% 54.4% 3.66
Seer [49] $\checkmark$ Cont. 94.4% 87.2% 79.9% 72.2% 64.3% 3.98
MoDE [23] $\checkmark$ Diffusion 96.2% 88.9% 81.1% 71.8% 63.5% 4.01
Seer-Large [49] $\checkmark$ Cont. 96.3% 91.6% 86.1% 80.3% 74.0% 4.28
DP3 [7] $\times$ Diffusion 28.7% 2.7% 0.0% 0.0% 0.0% 0.31
MDT [48] $\times$ Diffusion 63.1% 42.9% 24.7% 15.1% 9.1% 1.55
3DA [27] $\times$ Diffusion 92.2% 78.7% 63.9% 51.2% 41.2% 3.27
MoDE (scratch) [23] $\times$ Diffusion 91.5% 79.2% 67.3% 55.8% 45.3% 3.39
CLOVER [47] $\times$ Diffusion 96.0% 83.5% 70.8% 57.5% 45.4% 3.53
Seer (scratch) [49] $\times$ Cont. 93.0% 82.4% 72.3% 62.6% 53.3% 3.64
Seer-Large (scratch) [49] $\times$ Cont. 92.7% 84.6% 76.1% 68.9% 60.3% 3.83

### v2

@@ 1 Introduction
The advent of diffusion-based Imitation Learning (IL) has allowed robots to carry out complex, long-horizon tasks from raw image observations [chi2023diffusionpolicy, black2024pi0]. RGB images are a common observation modality for diffusion policies due to their ubiquitousness and rich semantic information. However, policies conditioned on only RGB images lack 3D geometric information about the scene. This 3D information is crucial for learning generalizable policies that can act precisely in complex 3D scenes, especially when using multiple camera views [jia2024towards, zhu2024point, peri2024pointcloudmodelsimprove, gyenes2024pointpatchrl]. An alternative modality is point clouds, unstructured sets of 3D points that preserve geometric shape, distances, and spatial relationships. In addition, points captured from multiple camera views can be transformed into a common reference frame and concatenated, yielding a natural and powerful way to fuse multiple cameras. Although numerous works use point clouds as an input modality [ze20243d, chisari2024learning, xian2023chaineddiffuser], their irregular structure limits the network architectures that can be used with them. In contrast, RGB images are on a regular grid and can be processed using convolutional operators, but are susceptible to changes in perspective and lighting.
@@
Current 3D processing approaches face fundamental limitations that create a critical gap between 3D geometric information and existing 2D vision architectures. Downsampling-based methods [ze20243d], as shown in Figure 1(a), suffer from an inherent information-fidelity tradeoff: they must dramatically reduce point density through techniques like Farthest Point Sampling (FPS)[qi2017pointnet++] to remain computationally tractable, inevitably discarding fine-grained geometric details essential for precise manipulation tasks. Feature-lifting approaches [ke20243d], as shown in Figure 1(b), face equally problematic limitations as they aggregate 2D features through depth averaging and 3D transformations, introducing information loss while struggling to maintain spatial structure during the lifting process. In this paper, we take inspiration from recent advances from the computer vision community in stereo reconstruction [wang2024DUSt3R, leroy2024MASt3R] to propose using point maps, as shown in Figure 1(c). Point maps encode 3D information in a regular, 2D grid of Cartesian coordinates. This results in a structured data type that can be used with standard architectures such as ResNet [he_deep_2016], ViT [dosovitskiy2021vit], or ConvNeXt [liu2022convnext]. This obviates the need for steps like K-Nearest Neighors (KNN) and Farthest Point Sampling (FPS) [qi2017pointnet++], which are computationally expensive operations common to point cloud methods [pang2022pointmae, chen_pointgpt_2023, gyenes2024pointpatchrl]. At the same time, because they are geometrically grounded, point maps from multiple views can be transformed into the same reference frame, increasing robustness to perturbations in camera perspective.
@@
We integrate point maps into a standard diffusion-based imitation learning framework based on EDM [karras2022elucidating] to demonstrate their effectiveness as a drop-in replacement for RGB images or point clouds. We validate the effectiveness of point map observations on two challenging benchmarks: RoboCasa [nasiriany2024robocasa] and CALVIN [mees2022calvin]. These benchmarks feature language-conditioned tasks and diverse scenes, and require spatial reasoning and long-term planning. Across both benchmarks, point map-based policies outperform baselines using RGB, depth maps, or point clouds, demonstrating superior learning efficiency and generalization. Our method is computationally efficient in training and inference, sometimes by an order of magnitude.
@@
Contributions: Our contributions are the following: 1) we propose PointMapPolicy (PMP), a method for diffusion-based imitation learning on point maps, a powerful observation modality that has never been used in diffusion imitation learning; 2) we achieve state of the art results among policies trained from scratch on the CALVIN benchmark [mees2022calvin], and outperform other observation modalities on RoboCasa [nasiriany2024robocasa]; 3) we present systematic ablations of point cloud processing methods, vision backbones (e.g. ResNet [he_deep_2016], ViT [dosovitskiy2021vit], ConvNeXt [liu2022convnext]), and paradigms for fusing color and geometry information.
@@ 2 Related Work
2D Visual Representations for Imitation Learning. Recent imitation learning approaches [chi2023diffusionpolicy, liu2025rdtb, kim2024openvla, reuss2025efficient, octo_2023] rely predominantly on 2D visual representations such as RGB images or videos. Such representations are widely utilized due to their capacity to capture rich textural and semantic information, as well as their accessibility through low-cost cameras. However, 2D image modalities have inherent limitations: they contain 3D information only implicitly, are vulnerable to viewpoint and lighting changes and occlusions, and typically underperform in tasks requiring detailed spatial reasoning and geometric alignment [jia2024towards, zhu2024point, peri2024pointcloudmodelsimprove, gyenes2024pointpatchrl].
@@
3D Visual Representations for Imitation Learning. To overcome these limitations, a growing amount of research incorporates explictly 3D representations such as depth maps, point clouds, or voxels. Voxel-based methods like C2F-ARM [coarse_to_fine_q_attention] and Perceiver-Actor [shridhar2022perceiveractor] voxelize point clouds and use a 3D-convolutional network for action prediction, but require high voxel resolution for precision tasks, resulting in high memory consumption and slow training. DP3 [ze20243d] encodes sparse point clouds using FPS, followed by a lightweight MLP to produce a compact embedding vector of the observation. While efficient, this approach discards local geometric structure that can be critical for fine-grained tasks. In contrast, 3D Diffuser Actor [ke20243d] computes tokens by lifting 2D image features into 3D space by using averaged depth information and camera parameters, and applies FPS after the first cross-attention layer. FPV-Net [donat2025towards] fuses RGB and point cloud modalities by injecting global and local image features into a point cloud encoder using adaptive normalization layers, but is still limited by the disadvantages of both modalities.
@@
Multi-View Representation. Complementary work, such as Robot Vision Transformer (RVT) [goyal2023rvt], avoids working directly with raw point clouds by proposing a novel multi-view representation. This approach re-renders the point cloud from a set of orthographic virtual cameras, deriving a 7-channel point map (RGBD + XYZ) from each view. RVT-2 [goyal2024rvt] improves this approach for high-precision tasks by introducing a multi-stage inference pipeline: it first identifies a region-of-interest, truncates the observation to this area of interest, and then runs policy inference. However, neither of these methods use action diffusion, instead relying on key-frame based manipulation with a motion planner [shridhar2022peract]. Furthermore, geometric and color information are fused naively at the channel level, whereas we investigate more sophisticated techniques for fusion.
@@
Diffusion-Policy Backbones. Due to the non-Markovian nature of human demonstrations, where successful decision-making often depends on histories of past observations and actions, early work used RNN-based architectures [mandlekar2021what], but struggled with vanishing gradients and limited scalability. This led to the adoption of Transformer-based architectures, which offer global attention and parallelism, enabling superior performance in tasks requiring long-horizon reasoning [shafiullah2022behavior, reuss_gcil, bharadhwaj2023roboagent], becoming the standard backbone for many methods [ze20243d, ke20243d, donat2025towards, liu2025rdtb]. However, Transformers are computationally intensive and scale quadratically with the sequence length, which limits the number of tokens that can be used to encode the observation.
@@
To mitigate these challenges, recent works [jia2024mailimprovingimitationlearning, jia2025xil] explore State Space Models (SSMs) like Mamba [gu2024mamba], achieving linear-time complexity and improved sample efficiency, particularly in low-data regimes. Additionally, recent recurrent architectures such as xLSTM [beck2024xlstm] provide an appealing balance, maintaining the temporal modeling strengths of traditional RNNs while introducing architectural innovations that improve gradient flow and expressiveness. Despite being less expressive than self-attention, xLSTM significantly reduces compute and memory costs, making it well-suited for real-time or resource-constrained applications. X-IL [jia2025xil] systematically compares different architectural parts, and finds that xLSTM performs competitively with Transformers in multi-modal imitation learning. Building on these insights, PointMapPolicy adopts xLSTM as its diffusion backbone, balancing temporal modeling capability with efficient training and inference.
@@ 3.1 Problem Formulation
Imitation Learning aims to learn a policy from expert demonstrations. Given a dataset of expert trajectories $\mathcal{D}_{\mathbf{\tau}}=\{\mathbf{\tau}_{i}\}_{i=1}^{N}$, where each trajectory $\mathbf{\tau}_{i}=\left((\mathbf{s}_{1},\mathbf{a}_{1}),(\mathbf{s}_{2},\mathbf{a}_{2}),\dots,(\mathbf{s}_{K},\mathbf{a}_{K})\right)$. The objective is to learn a policy $\pi(\bm{\bar{a}}|\mathbf{s})$ that maps observations $\mathbf{s}$ to a sequence of actions $\bm{\bar{a}}=(\mathbf{a}_{k},\mathbf{a}_{k+1},\dots,\mathbf{a}_{k+H})$. Predicting sequences of actions, i.e. action chunking, allows for more temporally consistent action prediction [zhao2023learning]. Each observation $\mathbf{s}$ contains multi-view RGB-D images and language instruction for the current trajectory.
@@ 3.2 Score-based Diffusion Policy
Our approach employs the EDM framework for continuous-time action diffusion [karras2022elucidating, reuss2023goal] to generate actions. Diffusion models are generative models that learn to generate new samples through learning to reverse a Gaussian Perturbation process. In PointMapPolicy, we apply a score-based diffusion model to formulate the policy representation $\pi_{\theta}(\bm{\bar{a}}|\mathbf{s})$. This perturbation and its inverse process can be expressed through a Stochastic Differential Equation (SDE):
@@
After training, we can generate new action sequences beginning with Gaussian noise by iteratively denoising the action sequence with a numerical Ordinary Differential Equation (ODE) solver. Our approach utilizes the DDIM-solver, a specialized numerical ODE-solver for diffusion models [song2021denoising] that enables efficient action denoising in a minimal number of steps. Across all experiments, our method uses $4$ denoising steps.
@@ 3.3 Observation Tokenization
We are given an observation $\mathbf{s}_{k}$ in step $k$ as well as a textual language instruction $z_{\text{lang}}$. The language instruction is first tokenized using a pretrained CLIP text encoder [radford2021learning] to generate language embeddings. For RGB inputs, we use Film-ResNet [perez2018film] with pretrained ImageNet weights to generate visual embeddings from the observation $\mathbf{s}_{k}$.
We define that a Point Map $X\in\mathbb{R}^{H\times W\times 3}$ is a dense 2D field of 3D points that establishes a one-to-one mapping between image pixels and 3D scene points. For an RGB image $I$ of resolution $H\times W$, the corresponding Point Map $X$ satisfies $I_{i,j}\leftrightarrow X_{i,j}$ for all pixel coordinates $(i,j)\in\{1\ldots H\}\times\{1\ldots W\}$, where each pixel intensity $I_{i,j}$ corresponds to a 3D point $X_{i,j}\in\mathbb{R}^{3}$ in world coordinates.
@@ 3.4 PointMapPolicy
Backbones. Given the multi-modal tokens from Section 3.3, a learnable positional embedding is added to each token. PMP uses a decoder-only backbone from X-IL [jia2025xil] with x-LSTM as the core computational unit. All tokens are concatenated as inputs to the X-Block, which is the diffusion score network $D_{\theta}$. While Transformers dominate most imitation learning policies, X-IL demonstrated that the recent recurrent architecture xLSTM excels in robot learning tasks. The core computational element within X-Block is the m-LSTM layer, which serves an analogous function to self-attention in Transformer architectures. The denoised action tokens produced by the X-Block are then used to guide the robot’s behavior, resulting in a policy that effectively leverages both the geometric precision of point maps and the rich semantic understanding from RGB images.
@@ 4 Simulation Experiments
Figure 4: Overview of Simulation and Real World Experiments used to test PointMapPolicy. From left to right: CALVIN Benchmark [mees2022calvin], RoboCasa [nasiriany2024robocasa], and our Real World Setup.
We conduct experiments on two simulation benchmarks RoboCasa [nasiriany2024robocasa] and CALVIN [mees2022calvin]. We aim to answer the following questions: Q1) How does PointMapPolicy compare to state-of-the-art 2D and 3D imitation learning policies? Q2) How do the fusion methods perform compared to other modalities? Q3) How does point map representation compare to other point cloud processing methods? Q4) Can current vision encoders effectively extract the geometric and semantic information from point maps required for robust decision-making?
RoboCasa: The RoboCasa benchmark [nasiriany2024robocasa] is a large-scale simulation framework designed to evaluate IL agents across a wide range of household manipulation tasks. Built on a physically realistic environment with rich visual rendering, RoboCasa supports task diversity, long-horizon behaviors, and fine-grained physical interactions, making it a compelling testbed for assessing both generalization and behavior diversity in policy learning. We use the RoboCasa benchmark to assess whether our proposed point map representation can enable effective learning and generalization across manipulation tasks of increasing complexity, object count, and behavioral variation.
CALVIN: The CALVIN benchmark [mees2022calvin] provides a large-scale framework for evaluating language-conditioned IL policies in visually rich, long-horizon manipulation tasks. The benchmark contains 34 distinct manipulation tasks such as button-pressing, drawer-opening, object-picking, and pushing. Each rollout consists of a sequence of 5 language instructions, and the agent must complete one task before proceeding to the next. Policies are evaluated on 1,000 such instruction chains per seed, and success is measured by the average number of correctly completed tasks in each sequence.
@@
Table 1: Success rate (%) for each task in RoboCasa [nasiriany2024robocasa]. The models were trained for 50 epochs with 50 human demonstrations per task and evaluated with 50 episodes for each task. The bold numbers highlight the best achieved success rate for that task among all the models.
Baselines: For RoboCasa, we benchmark against Behavioral Cloning (BC) [nasiriany2024robocasa], GR00T-N1 [bjorck2025gr00t], 3D Diffusion Policy (DP3) [ze20243d], and 3D Diffuser Actor (3DA) [ke20243d]. Note that GR00T-N1 results use $100$ demonstrations, while all other methods use $50$ human demonstrations. To systematically evaluate the effectiveness of our representation, we further compare other against image-based baselines using only RGB data (RGB), and only depth data (Depth). We then compare PMP against multiple variants introduced in Section 3.4: PMP-6ch directly uses 6-channel point maps as inputs, and PMP-xyz only uses xyz coordinates as inputs. All five methods share the same architectures and parameters for fair comparison. Details can be found in Appendix 6.
For CALVIN, we primarily compare our approach against models without robot-specific pretraining, though we include all results for reference. DP3, 3DA, and CLOVER [bu2024clover] are selected as representative policies using RGB-D inputs. MDT [reuss2024multimodal], MoDE [reuss2025efficient], and Seer [zhou2024seer] are selected as RGB-based policies. Further details of these baselines can be found in Appendix A.2.
@@
SuSIE [black2023zero] $\checkmark$ Diffusion 87.0% 69.0% 49.0% 38.0% 26.0% 2.69
GR-1 [wu2023unleashing] $\checkmark$ Cont. 85.4% 71.2% 59.6% 49.7% 40.1% 3.06
OpenVLA [kim2024openvla] $\checkmark$ Discrete 91.3% 77.8% 62.0% 52.1% 43.5% 3.27
RoboDual [bu2024towards] $\checkmark$ Diffusion 94.4% 82.7% 72.1% 62.4% 54.4% 3.66
Seer [zhou2024seer] $\checkmark$ Cont. 94.4% 87.2% 79.9% 72.2% 64.3% 3.98
MoDE [reuss2025efficient] $\checkmark$ Diffusion 96.2% 88.9% 81.1% 71.8% 63.5% 4.01
Seer-Large [zhou2024seer] $\checkmark$ Cont. 96.3% 91.6% 86.1% 80.3% 74.0% 4.28
DP3 [ze20243d] $\times$ Diffusion 28.7% 2.7% 0.0% 0.0% 0.0% 0.31
MDT [reuss2024multimodal] $\times$ Diffusion 63.1% 42.9% 24.7% 15.1% 9.1% 1.55
3DA [ke20243d] $\times$ Diffusion 92.2% 78.7% 63.9% 51.2% 41.2% 3.27
MoDE (scratch) [reuss2025efficient] $\times$ Diffusion 91.5% 79.2% 67.3% 55.8% 45.3% 3.39
CLOVER [bu2024clover] $\times$ Diffusion 96.0% 83.5% 70.8% 57.5% 45.4% 3.53
Seer (scratch) [zhou2024seer] $\times$ Cont. 93.0% 82.4% 72.3% 62.6% 53.3% 3.64
Seer-Large (scratch) [zhou2024seer] $\times$ Cont. 92.7% 84.6% 76.1% 68.9% 60.3% 3.83